OpenStudy (anonymous):

Hi everyone! Can someone help me factor r^3 + 3r^2 - 4 = 0 please? My book says it is obvious that by inspection one of the roots are r=1...not so obvious to me. Thanks! :o)

OpenStudy (anonymous):

\[1+3-4=0\] is more or less obvious

OpenStudy (anonymous):

if \(r=1\) you add the coefficients that is always true

OpenStudy (welshfella):

well if you plug in 1 for r you get 1^3 + 3*1^2 - 4 = 0

OpenStudy (anonymous):

@satellite...why is it okay to do 1+3-4=0 i mean what about grouping or all the other methods of solving...please explain why it's okay to just look at the coefficients in that way

OpenStudy (misty1212):

HI!! why does everyone start with "well" seem rather affected to me as does "you just..."

OpenStudy (anonymous):

1 to any power is 1

OpenStudy (anonymous):

so for example if you have \[f(x)=4x^3+5x^2+2x-1\] then \[f(1)=4+5+2-1\]

OpenStudy (ribhu):